OpenStudy (anonymous):

how can I solve for x in terms of other variables? Example included: p-6/7x=2q

OpenStudy (anonymous):

so to solve this, I want to get rid of any fractions, because fractions are evil. so i would multiply both sides by 7x. Next i would want to get x on its own. so i would divide both sides by (2 x 7 x q)

OpenStudy (anonymous):

I don't quite understand

OpenStudy (anonymous):

okay multiply both sides of your equation by 7x and type what you get

OpenStudy (anonymous):

oh I made a typo... the fraction in the equation is -6/7

OpenStudy (anonymous):

Would that mean I also multiply by -7, right?

OpenStudy (anonymous):

\[\frac{-6}{7x} = 2q\] ?

OpenStudy (anonymous):

is that the equation

OpenStudy (anonymous):

no, the equation is p-6/7x=2q

OpenStudy (anonymous):

\[\frac{p-6}{7x} = 2q\] ?

OpenStudy (anonymous):

\[p - \frac{6}{7x} = 2q\] ?

OpenStudy (anonymous):

use brackets in your equation

OpenStudy (anonymous):

\[p-\frac{ -6 }{ 7x }=2q\]

OpenStudy (anonymous):

so that last one with brackets would be p -(6/7x) = 2q

OpenStudy (anonymous):

p minus minus 6, or p minus 6?

OpenStudy (anonymous):

first one?

OpenStudy (anonymous):

yeah p- negative six over seven and the x is in the middle of the equation next to the line between the 6 and 7

OpenStudy (anonymous):

cool awesome

OpenStudy (anonymous):

okay so - - 6 = + 6 right?

OpenStudy (anonymous):

opps....forget the p minues negative six its just p minus 6

OpenStudy (anonymous):

so....p minues negative 6 over 7x equals 2q

OpenStudy (anonymous):

right okay so firstly you want to minus both sides by p. type what you get

OpenStudy (anonymous):

i would get -6/7x=2q-p

OpenStudy (anonymous):

yeah you would, now multiply both sides by 7x, write the right hand side as 7x(2q-p)

OpenStudy (anonymous):

which gives you -6 = 7x(2q - p) right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

cool, now divide both sides by 7(2q - p)

OpenStudy (anonymous):

wait...now i use distributive property...right?

OpenStudy (anonymous):

no keep it inn the brackets its easier

OpenStudy (anonymous):

but, them I would divide -6 by 7(2q-p)

OpenStudy (anonymous):

yeah exactly

OpenStudy (anonymous):

and you would have x on its own in terms of the variables q and p

OpenStudy (anonymous):

I don't quite get how I would actually divide it by that.

OpenStudy (anonymous):

you dont, you just leave it in that form \[\frac{-6}{7(2q-p)} = x\]

OpenStudy (anonymous):

oh....Can you help me in the other ones? I thought i'd get the hang of it with one

OpenStudy (anonymous):

sure post anotehr here

OpenStudy (anonymous):

another*

OpenStudy (anonymous):

alright....\[\frac{ 2 }{ 3 }ab-\frac{ 1 }{ 5 }x=2\]

OpenStudy (anonymous):

ugh thats an ugly oen, but suprisingly easy

OpenStudy (anonymous):

so first step is to try and get as little stuff on the side of x as possible. is there anything you do to get that?

OpenStudy (anonymous):

I multiply by 5?

OpenStudy (anonymous):

i would \[-\frac{2}{3}ab\] first.

OpenStudy (anonymous):

but multiplying everything by 5 will also work, take your pick

OpenStudy (anonymous):

so multiply by -3?

OpenStudy (anonymous):

oh alright I'll multiply by 5

OpenStudy (anonymous):

okay type what you get

OpenStudy (anonymous):

I get \[\frac{ 10 }{ 6 }ab-1x=10\]

OpenStudy (anonymous):

awesome, any idea how you can get just -1x on the left side?

OpenStudy (anonymous):

in one move :)

OpenStudy (anonymous):

add x so then you would have \[\frac{ 10 }{ 6 }ab =10+x\]

OpenStudy (anonymous):

yes that will be just fine. now what can you do to get x on its own

OpenStudy (anonymous):

then....multiply each side by 6 so then I would get: \[10ab=1000+x\]

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

well...to get x on its own, i would have to subtract 10

OpenStudy (anonymous):

sorry, got ahead of myself

OpenStudy (anonymous):

yes exactly! and what would that give you

OpenStudy (anonymous):

if you subtracted 10

OpenStudy (anonymous):

i would get....hmmm how do I deal with the equation?

OpenStudy (anonymous):

what do you mean? i think from here: \[\frac{ 10 }{ 6 }ab =10+x\] by subtracting 10: \[\frac{ 10 }{ 6 }ab - 10= x\]

OpenStudy (anonymous):

agree?

OpenStudy (anonymous):

oh! ok

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

I would appreciate trying two more but let me do at least one on my own

OpenStudy (anonymous):

well just try and if you get stuck message me

OpenStudy (anonymous):

\[\frac{ 5 }{ 2 }(h-x)=\frac{ k }{ 4 }\]

OpenStudy (anonymous):

I would multiply by 2 each side

OpenStudy (anonymous):

thats a good idea

OpenStudy (anonymous):

so then i get....\[5(2h-2x)=\frac{ 2k }{ 8 }\]

OpenStudy (anonymous):

yeah. now get rid of that 5 to free the x from the bracket

OpenStudy (anonymous):

that would create.....\[2h-2x=\frac{ 2k }{ 8 }-5\]

OpenStudy (anonymous):

okay now what do you think you should do? remember your trying to get as little stuff on the side of x as possible

OpenStudy (anonymous):

I would subtract 2h which would create....\[-2x=\frac{ 2k }{ 8 }-5-2h\]

OpenStudy (anonymous):

yeah! now what do you think?

OpenStudy (anonymous):

x doesn't need to be divided, right? that would be the answer?

OpenStudy (anonymous):

well, we want to get rid of the 2 next to x, forget about the - sign for now

OpenStudy (anonymous):

well dividing by -2 would fix it

OpenStudy (anonymous):

sure it would. so divide everything by -2 and you have done it

OpenStudy (anonymous):

your good at these

OpenStudy (anonymous):

well probably not as fluent as you since i just learned it

OpenStudy (anonymous):

it just becomes second nature when you start learning about other things. you will start taking it for granted.

OpenStudy (anonymous):

well...I've got other homework to do. Thanks.